Grade 12General Physics

three blocks of mass m1=2 kg, m2 =4kg and m3 =6kg are connected by light string on a frictionless inclined plane of 60°. A force F of 120 N is applied upwards along the incline to the uppermost block, causing an upward movement of the block. calculate the acceleration of the blocks and tension in the strings.

rudra pratap
8 Years agoGrade 12

1 Answer

jitender
8 Years ago
F - T1-m1gsin60 = m1a_______ 1st eqnT1-T2-m2gsin60= m2a_______2nd eqnT3 -m3gsin60= m3a_____3rd.eqnAnd F-(m1+m2+m3)gsin60 = (m1+m2+m3)aFirst find a then T3 from 3rd then T2 from 2nd and then T1 from 1st eqn